OpenStudy (anonymous):

what are the solutions of the system? y=x^2-x+2 y=x+5

OpenStudy (anonymous):

set those two equation equal to each other in y:

OpenStudy (anonymous):

x^2 - x + 2 = x + 5 x^2 - 2x - 3 = 0 can you take it from here?

OpenStudy (anonymous):

x^2 -x+2=x+5 and solve it

OpenStudy (anonymous):

oh dpalnc did the same :)

OpenStudy (anonymous):

i dont think so..

OpenStudy (anonymous):

all yours @kittu ... :)

OpenStudy (anonymous):

@Mujod122 i didn't get you @dpaInc :)

OpenStudy (anonymous):

i have no idea how to do this

OpenStudy (anonymous):

oh yeah they don't have real roots

OpenStudy (anonymous):

sorry i should've checked

OpenStudy (anonymous):

oh no they have real roots

OpenStudy (anonymous):

x^2 - 2x - 3 = 0 and the roots are -1 and +3

OpenStudy (anonymous):

now substitute these two in the equation and find the value of y

OpenStudy (anonymous):

x=3,-1 ??

OpenStudy (anonymous):

yeah man.... what is the problem you have then ? u get two answers for this because it is a quadratic equation

OpenStudy (anonymous):

idk i got confused thanks

OpenStudy (anonymous):

I can help you if you tell me the part you don't understand :)